When painting the walls of an art museum, the painters found that it took 5 men to paint 4 walls in 6 hours. Then, 9 more men were added to the job. How many hours would it take them to do 28 rooms?
Answer by lwsshak3(11628)

When painting the walls of an art museum, the painters found that it took 5 men to paint 4 walls in 6 hours. Then, 9 more men were added to the job. How many hours would it take them to do 28 rooms?
**
It took 5*6=30 man-hours to complete painting 4 walls or one room
28 rooms will require 30*28=840 man hours
men on job=5+9=13
Hours to paint 28 rooms=840/13≈64.62 hrs